### v1

@@ 3.4 Computational efficiency
In this section, we analyse the computational efficiency of our proposed shadow removal method in comparison with several state-of-the-art models. The comparison is based on key metrics including frames per second (FPS), inference time, computational complexity in giga FLOPs (FLOPs), and the number of parameters (Params). All models were tested on the ISTD dataset using a 2080Ti GPU device. As shown in Table 5 and Fig. 2, our method significantly outperforms the other models in terms of FPS, achieving a rate of 66. This is considerably higher than the next best model, SP+M-Net, which achieves an FPS of 10. Our method also boasts the shortest inference time of 0.015 s, demonstrating its efficiency. In terms of computational complexity, our method requires only 7.5 giga FLOPs, which is significantly lower than the other models. For instance, the DHAN model requires 262.9 giga FLOPs, which is over 35 times more than our method. Furthermore, our method uses only 16.1M parameters, which is less than most of the other models. The ShadowFormer model uses slightly fewer parameters (9.3M), but it achieves a lower FPS of 8 compared to our method’s 66. In conclusion, our method not only delivers superior performance but also demonstrates impressive computational efficiency, making it a highly effective solution for shadow removal tasks for field robotics.
Table 5: Performance and Complexity Analysis of Shadow Removal Methods: A comparison of our method with state-of-the-art models on key metrics including PSNR performance, inference time, frames per second (FPS), number of parameters (Params), and computational complexity (FLOPs). All models were tested on the ISTD dataset using a 2080Ti GPU device.
Method FPS Time FLOPs PSNR Params DHAN (Cun et al., 2020) 9 0.117 262.9 29.12 21.8 SP+M-Net (Le and Samaras, 2019) 10 0.105 160.1 25.08 141.2 G2R (Liu et al., 2021) 4 0.254 113.9 25.07 113.9 Fu et al. (2021) 7 0.155 160.3 27.19 143.0 ShadowFormer (Guo et al., 2023) 8 0.127 100.9 32.21 9.3 Ours 66 0.015 7.5 32.76 16.1
@@
3.5 Application of Shadow Removal in Field Robotics
Our shadow removal method presents substantial benefits for two critical areas of field robotics: real-time image processing and image processing during dataset labelling.
3.5.1 Real-time Image Processing
Field robots often operate in dynamic environments with varying lighting conditions. This can significantly impact the uniformity of captured images, hindering the performance of computer vision pipelines within the robot. Our shadow removal method addresses this challenge by enabling real-time image processing. By removing shadows from the captured frames, the method ensures consistent image quality and improves the clarity of crucial features for object detection and recognition tasks. This is particularly valuable in agricultural settings where environmental lighting fluctuates throughout the day. Consistent image quality allows the robot’s computer vision system to perform more accurate object detection and recognition tasks, leading to improved decision-making and overall performance.
3.5.2 Image Processing during Dataset Labelling
@@
Figure 9 presents real-world images captured by a weed control robot, before and after shadow removal. The clear improvement in image quality and visibility of the weeds demonstrates the effectiveness of our method in improving the labelling process.
Tables 6 and 7 present the measured inference times (ms) and frame rates (fps) of the shadow algorithm, respectively. These tables quantify the performance of the algorithm on a Jetson-Orin AGX platform in various configurations. They explore the impact of input size (256x256 and 512x512 pixels) and weight precision (float32, float16, int8) for both PyTorch and TensorRT implementations. Analyzing these results is crucial for selecting the optimal configuration for real-world applications, striking a balance between speed and accuracy requirements.
@@
Table 6: Inference times (ms) of the shadow algorithm measured on a Jetson Orin AGX embedded GPU platform varying input size (512$\times$512 and 256$\times$256 px) and weight precision (float32, float16, int8) for PyTorch and TensorRT implementations.
@@
Table 7: Frame rates (fps) of the shadow algorithm measured on a Jetson Orin AGX embedded GPU platform varying input size (512$\times$512 and 256$\times$256 px) and weight precision (float32, float16, int8) for PyTorch and TensorRT implementations.
@@ 3.6 Ablation Study
An ablation study was conducted to understand the contribution of each component of the loss function to the performance of the ShadowRemovalNet model. The loss function is composed of three main components: the enhancement loss $L_{e}$, the KL divergence $(L_{m}+L_{s})$, and the boundary loss $L_{b}$. Each of these components is weighted by a corresponding weight parameter ($\alpha$, $\beta$, and $\gamma$ respectively) in the total loss function:
$L=\alpha L_{e}+\beta(L_{m}+L_{s})+\gamma L_{b}$
In the ablation study, each component of the loss function was removed one at a time, and the performance of the model was evaluated without that component. This allowed us to quantify the impact of each component on the overall performance of the model. The results of the ablation study, Tab. 8, showed that each component of the loss function contributes significantly to the performance of the ShadowRemovalNet model. The enhancement loss $L_{e}$ was found to be crucial for generating high-quality shadow-free images, while the KL divergence $(L_{m}+L_{s})$ played a key role in aligning the posterior and prior distributions in the latent space. The boundary loss $L_{b}$ was particularly important for enhancing the supervision of pixels near the shadow boundaries, leading to more accurate shadow removal in these areas.
Table 8: Contributions of the components of the loss function to the performance of the ShadowRemovalNet model tested on ISTD (Wang et al., 2018) datasets. .
@@ 4 Discussion
Implications. The findings from our study have significant implications for field robotics. Our method for shadow removal using unpaired data demonstrates that it is possible to effectively identify and remove shadows from images, even without a direct comparison to shadow-free ground truth images. This opens up new possibilities for improving the accuracy and efficiency of computer vision in various outdoor lighting conditions. The use of unpaired data also means that our method can be applied in a wider range of scenarios, as it does not require the collection of paired shadow and shadow-free images. This could potentially simplify the data collection process and make our method more accessible for use in different environments and lighting conditions. Moreover, our method’s ability to handle complex and diverse image data suggests that it could be applied in a wide variety of field robotics applications, where challenges with lighting and shadows may exist.
@@
Relationship with Unsupervised Method. Shadow image degradation is complex and irreversible, making it infeasible to directly supervise a shadow removal model with ground truth data. Therefore, unsupervised methods are better suited for this task. Rather than developing a new unsupervised approach, our work focuses on effectively utilizing synthetic labels from existing shadow removal datasets. While not perfect, these synthetic reference maps of shadow-free images can be easily obtained. We propose a framework that leverages conditional variational autoencoders and our proposed module to learn an enhancement model from these imperfect references. In summary, both our approach and unsupervised methods aim to train a shadow removal model without ground truth data. However, our framework provides a novel perspective, i.e., transforming the problem into distribution estimation and consensus building using readily available, but biased references. This represents a distinct approach from existing unsupervised techniques.
@@
Training Dataset and Model Development. The development of robust shadow removal algorithms is significantly influenced by the quality and quantity of available training data. Currently, the scarcity of high-quality image pairs, each consisting of a shadow and its corresponding shadow-free version of the same scene, poses a considerable challenge. The largest available datasets for shadow removal, such as the state-of-the-art SRD and ISTD datasets, contain only around 100 unique scenes. This limited dataset size may not capture the diversity and complexity of real-world scenes, leading to suboptimal generalization to unseen scenes and inferior performance in shadow removal tasks. Models trained on such limited datasets may struggle to understand the variations in lighting conditions, scene composition, and other factors that affect the appearance of shadows. This could result in a failure to maintain consistent colors between the shadow and shadow-free regions, leading to unnatural or inconsistent results in shadow removal. The difficulty in collecting a large dataset with paired shadow/shadow-free images, given the various types of shadow masks and scenes, further compounds these challenges. To overcome these limitations, we have collected a substantial dataset of 10,000 natural shadow-free images from the internet. This large-scale collection significantly enhances the robustness and generalization ability of our shadow removal method. By synthesizing shadow images from this extensive dataset, we can train our shadow removal algorithms more effectively, thereby achieving superior performance in shadow removal tasks. This approach demonstrates the potential of using internet-sourced, shadow-free images to improve the training of shadow removal algorithms.
@@
Limitations. Despite its strengths, our method also has some limitations. One of the main limitations is that it may struggle with extremely complex shadows or lighting conditions. While our method performs well under a variety of conditions, there may be certain scenarios where the complexity of the shadows exceeds the capabilities of our method. Another limitation is the reliance on unpaired data. While this allows for greater flexibility in data collection, it also means that our method does not have a direct comparison to a shadow-free ground truth image. This could potentially impact the accuracy of the shadow removal process. Finally, while our method has demonstrated strong performance on the datasets we tested, its performance on other datasets or real-world scenarios remains to be seen. It is possible that our method may not perform as well on different types of images or in different environments.
@@ 5 Conclusions
Our study presents a novel method for shadow removal using unpaired data. Our method effectively handles the inherent complexity of shadow removal, delivering high-quality results even in challenging lighting conditions. Despite the complexity of the task, our method demonstrates impressive computational efficiency, outperforming several state-of-the-art models in terms of key performance metrics. However, like any method, ours is not without its limitations. It may struggle with extremely complex shadows or lighting conditions, and its performance on other datasets or real-world scenarios remains to be seen. Despite these limitations, our method opens up new possibilities for improving the accuracy and efficiency of outdoor computer vision in various lighting conditions. Looking ahead, there are several potential avenues for future research. These include improving the handling of complex shadows, testing our method on a wider range of datasets, optimizing computational efficiency, and exploring applications in other related fields. Through continued research and development, we believe our method has the potential to make significant contributions to computer vision for field robotics.
@@ 5.1 CO2 Emission Related to Experiments
[…]
Data Availability
Data will be made available on request
@@ Acknowledgement
Additional Information
Competing interests The authors declare no competing interests.
[…]

### v2

@@ 5.4 Computational efficiency
In this section, we analyse the computational efficiency of our proposed shadow removal method in comparison with several state-of-the-art models. FieldNet achieves 66 FPS on an Nvidia 2080Ti through its lightweight U-Net-based architecture (2.7M parameters, 7.37G FLOPs) and single-pass inference, avoiding the iterative processes of GANs (e.g., DHAN: 9 FPS, 262.9G FLOPs) or diffusion models (e.g., Mei et al., 2024:  0.5-1 FPS). Compared to ShadowFormer (8 FPS, 100.9G FLOPs) and ST-CGAN (slow due to 29.2M parameters), FieldNet offers 7-9x higher speed with significantly lower computational demands, as shown in Tab. VI and Fig. 2, making it highly practical for real-time field robotics.
TABLE VI: A comparison of our method with state-of-the-art models on key performance and complexity analysis metrics including PSNR, inference time, frames per second (FPS), number of parameters (Params), and computational complexity (gigaFLOPs). All models were tested on the ISTD dataset using a 2080Ti GPU device.
Method FPS Time FLOPs PSNR Params DHAN (Cun et al., , 2020) 9 0.117 262.9 29.12 21.8 SP+M-Net (Le and Samaras, , 2019) 10 0.105 160.1 25.08 141.2 G2R (Liu et al., , 2021) 4 0.254 113.9 25.07 113.9 Fu et al., (2021) 7 0.155 160.3 27.19 143.0 BMnet (Guo et al., , 2023) 16 0.062 11.0 30.28 0.4 ShadowFormer (Guo et al., , 2023) 8 0.127 100.9 32.21 9.3 FieldNet (Ours) 66 0.015 7.5 32.76 2.7
@@
5.5 Real-world case study of FieldNet in field robotics
In this section, we present the results of applying ShaodowRemovalNet to a real-world field robotic setting. We show that FieldNet presents substantial benefits for two critical areas of field robotics: real-time shadow removal and image enhancement during dataset labelling.
5.5.1 Real-time shadow removal in agriculture field robotic
Field robots often operate in dynamic environments with varying lighting conditions. This can significantly impact the uniformity of captured images, hindering the performance of computer vision pipelines within the robot. FieldNet addresses this challenge by enabling real-time image processing. By removing shadows from the captured frames, the method ensures consistent image quality and improves the clarity of crucial features for object detection and recognition tasks. This is particularly valuable in agricultural settings where environmental lighting fluctuates throughout the day. Consistent image quality allows the robot’s computer vision system to perform more accurate object detection and recognition tasks, leading to improved decision-making and overall performance.
Figure 9 presents real-world images captured by a weed control robot (Azghadi et al., , 2024), before and after shadow removal. The clear improvement in image quality and visibility of the weeds demonstrates the effectiveness of our method in removing shadows.
In addition, Tables VII and VIII present the measured inference times (ms) and frame rates (fps) of the shadow algorithm, respectively. These tables quantify the performance of the algorithm on a Jetson-Orin AGX platform in various configurations. They explore the impact of input size (256x256 and 512x512 pixels) and weight precision (float32, float16, int8) for both PyTorch and TensorRT implementations. Analyzing these results is crucial for selecting the optimal configuration for real-world applications, striking a balance between speed and accuracy requirements.
5.5.2 Image enhancement for improved dataset labelling
@@
As shown in Figure 9, the image quality and visibility of the weeds achieved after applying FieldNet signifies the effectiveness of our method in improving the labelling process, which could lead to improved deep learning of weed images.
Figure 10: Qualitative analysis of FieldNet’s sub-optimal performance. Examples include (top row) high-contrast shadow edges and uneven illumination and (bottom row) complex textures where residual artefacts remain after shadow removal. These edge cases highlight the challenges of achieving consistent shadow-free images in diverse scenarios.
TABLE VII: Inference times (ms) of the shadow algorithm measured on a Jetson Orin AGX embedded GPU platform varying input size (512$\times$512 and 256$\times$256 px) and weight precision (float32, float16, int8) for PyTorch and TensorRT implementations.
@@
TABLE VIII: Frame rates (fps) of the shadow algorithm measured on a Jetson Orin AGX embedded GPU platform varying input size (512$\times$512 and 256$\times$256 px) and weight precision (float32, float16, int8) for PyTorch and TensorRT implementations.
@@
5.6 Qualitative Analysis of Sub-Optimal Performance
While FieldNet demonstrates robust shadow removal in most scenarios, there are edge cases where its performance is less effective. Figure 10 illustrates such examples, including high-contrast shadow edges, complex textures, and uneven illumination. In cases with high-contrast edges (top row), residual artifacts persist due to challenges in precisely delineating sharp boundaries, reducing PSNR by approximately 1-2 dB in these regions (Tab. IX). Complex textures (bottom row) confuse feature modulation, leading to inconsistent color correction and minor artifacts. These limitations suggest areas for future improvement, such as refining the Probabilistic Enhancement Module (PEM) to better handle intricate shadow patterns or incorporating domain-specific augmentations during training to enhance texture robustness, as discussed in Sec. 6.
@@ 5.7 Ablation Study
An ablation study was conducted to understand the contribution of each component of the loss function and the Physics-Guided Enhancement Module (PEM) to the performance of the FieldNet model. The loss function is composed of three main components: the enhancement loss $L_{e}$, the KL divergence $(L_{m}+L_{s})$, and the boundary loss $L_{b}$. Each of these components is weighted by a corresponding weight parameter ($\alpha$, $\beta$, and $\gamma$ respectively) in the total loss function presented in Eq. 17.
To evaluate the impact of PEM, we generated 10,000 synthetic samples using the shadow synthesis process described in Section 3.2. These samples were used to train and validate the model, ensuring a robust evaluation of PEM’s contribution. The results, as shown in Tab. IX, demonstrate that PEM significantly improves the model’s ability to handle complex lighting conditions and generate high-quality shadow-free images.
In the ablation study, each component of the loss function was removed one at a time, and the performance of the model was evaluated without that component. This allowed us to quantify the impact of each component on the overall performance of the model. The results of the ablation study, as shown in Tab. IX, revealed the following:
  • The enhancement loss $L_{e}$ was found to be crucial for generating high-quality shadow-free images, as it directly optimizes the pixel-wise reconstruction accuracy.
  • The KL divergence $(L_{m}+L_{s})$ played a key role in aligning the posterior and prior distributions in the latent space, enabling the model to incorporate diversity into the shadow removal process.
  • The boundary loss $L_{b}$ was particularly important for enhancing the supervision of pixels near the shadow boundaries, leading to more accurate shadow removal in these areas.
TABLE IX: Contributions of the components of the loss function and PEM to the performance of the FieldNet model tested on the ISTD (Wang et al., , 2018) dataset.
@@
The ablation study highlights the importance of each component in the loss function and the significant contribution of PEM to the overall performance of FieldNet. Specifically, the novel loss function increases PSNR by 1.3 dB (from 36.54 to 37.84) by improving boundary accuracy through edge-aware supervision. Mask dissociation enhances SSIM by 0.02 (from 0.944 to 0.986) by refining shadow edges via body and detail mask separation. PEM boosts PSNR by 0.8 dB (from 36.88 to 37.84) by diversifying shadow-free outputs, enhancing robustness to lighting variations. Together, these components elevate FieldNet’s performance to a PSNR of 30.88, SSIM of 0.964, and RMSE of 4.57 on ISTD, far surpassing the baseline (PSNR: 20.76), as evidenced in Tab. IX. This synergy ensures high-quality shadow removal across diverse scenarios.
@@ 6 Discussion
Implications. The findings from our study have significant implications for real-time shadow removal applications such as field robotics. Our method for shadow removal using unpaired data demonstrates that it is possible to effectively identify and remove shadows from images, even without a direct comparison to shadow-free ground truth images. This opens up new possibilities for improving the accuracy and efficiency of computer vision in various outdoor lighting conditions. The use of unpaired data also means that our method can be applied in a wider range of scenarios, as it does not require the collection of paired shadow and shadow-free images. This could potentially simplify the data collection process and make our method more accessible for use in different environments and lighting conditions. Moreover, our method’s ability to handle complex and diverse image data suggests that it could be applied in a wide variety of field robotics applications, where challenges with lighting and shadows may exist. We demonstrated this by applying FieldNet to a real-world precision agriculture robotic application for weed detection.
@@
Relationship with unsupervised methods. Shadow image degradation is complex and irreversible, making it infeasible to directly supervise a shadow removal model with ground truth data. Therefore, unsupervised methods are better suited for this task. Rather than developing a new unsupervised approach, our work focuses on effectively utilizing synthetic labels from existing shadow removal datasets. While not perfect, these synthetic reference maps of shadow-free images can be easily obtained. We propose a framework that leverages conditional variational autoencoders and our proposed module to learn an enhancement model from these imperfect references. In summary, both our approach and unsupervised methods aim to train a shadow removal model without ground truth data. However, our framework provides a novel perspective, i.e., transforming the problem into distribution estimation and consensus building using readily available, but biased references. This represents a distinct approach from existing unsupervised techniques.
@@
Training dataset and model development. The development of robust shadow removal algorithms is significantly influenced by the quality and quantity of available training data. Currently, the scarcity of high-quality image pairs, each consisting of a shadow and its corresponding shadow-free version of the same scene, poses a considerable challenge. The largest available datasets for shadow removal, such as the state-of-the-art SRD and ISTD datasets, contain only around 100 unique scenes. This limited dataset size may not capture the diversity and complexity of real-world scenes, leading to suboptimal generalization to unseen scenes and inferior performance in shadow removal tasks. Models trained on such limited datasets may struggle to understand the variations in lighting conditions, scene composition, and other factors that affect the appearance of shadows. This could fail to maintain consistent colours between the shadow and shadow-free regions, leading to unnatural or inconsistent results in shadow removal. The difficulty in collecting a large dataset with paired shadow/shadow-free images, given the various types of shadow masks and scenes, further compounds these challenges.
To overcome these limitations, we collected a substantial dataset of 10,000 natural shadow-free images from the internet. This large-scale collection significantly enhances the robustness and generalization ability of our shadow removal method. By synthesizing shadow images from this extensive dataset, we can train our shadow removal algorithms more effectively, thereby achieving superior performance in shadow removal tasks. This approach demonstrates the potential of using internet-sourced, shadow-free images to improve the training of shadow removal algorithms.
@@
FieldNet’s efficiency stands out when compared to transformer-based methods like ShadowFormer (Guo et al., , 2023) and diffusion models (Mei et al., , 2024). Its real-time performance (66 FPS) surpasses ShadowFormer’s 8 FPS and diffusion models’ 0.5-1 FPS, driven by a lightweight architecture (2.7M parameters vs. 9.3M for ShadowFormer), making it highly suitable for field robotics where computational resources are limited. Additionally, its robust performance across ISTD, ISTD+, SRD, and SBU-Timelapse datasets highlights strong generalization, supported by unpaired training and a large synthetic dataset.
Limitations. Despite its strengths, our method also has some limitations. One of the main limitations is that it may struggle with extremely complex shadows or lighting conditions. While our method performs well under a variety of conditions, there may be certain scenarios where the complexity of the shadows exceeds the capabilities of our method. Another limitation is the reliance on unpaired data. While this allows for greater flexibility in data collection, it also means that our method does not have a direct comparison to a shadow-free ground truth image. This could potentially impact the accuracy of the shadow removal process. Finally, while our method has demonstrated strong performance on the datasets we tested and a real-world agriculture task, its performance on other datasets or real-world scenarios remains to be seen. It is possible that our method may not perform as well on different types of images or in different environments, as this is the case for any other shadow removal methods.
Despite its strengths, FieldNet exhibits limitations in handling high-contrast shadow edges and complex textures, where residual artifacts reduce performance by 1-2 dB in PSNR (Tab. IX, Fig. 10). Compared to diffusion models, which may better generalize to extreme lighting due to iterative refinement, FieldNet’s single-pass approach may miss fine details. Generalization could be further improved by incorporating real-world paired data and multi-domain augmentations, as its current reliance on synthetic shadows limits adaptability to rare scenarios.
@@ 7 Conclusions
In this study, we introduced FieldNet, a novel method for efficient and robust shadow removal using unpaired data. FieldNet addresses key challenges in shadow removal, including computational efficiency, boundary artefacts, and generalisation to diverse lighting conditions. The proposed approach incorporates a novel loss function that enhances supervision at shadow boundaries, significantly reducing errors compared to state-of-the-art methods. Additionally, the PEM ensures adaptive shadow removal without sacrificing real-time performance.
Our experimental results demonstrate FieldNet’s superior performance across multiple benchmarks, achieving a processing speed of 66 FPS on Nvidia 2080Ti hardware while maintaining high accuracy in terms of structural and perceptual quality metrics. The method’s scalability and robustness were further validated through its application in a real-world field robotics scenario for precision agriculture. Specifically, FieldNet improves robotic systems in two impactful ways: (1) enabling real-time shadow removal to enhance image clarity for downstream tasks, and (2) improving human labelling accuracy during dataset preparation for deep learning model training.
Overall, FieldNet opens new avenues for advancing computer vision in outdoor environments, particularly in resource-constrained applications like autonomous farming and robotics. The method’s lightweight architecture and adaptability make it a promising tool for enhancing the accuracy and efficiency of vision-based systems operating under diverse and challenging lighting conditions.
@@ 7.1 CO2 Emission Related to Experiments
Experiments were conducted using a private infrastructure, which has a carbon efficiency of 0.432 kgCO2eq/kWh. A cumulative of 850 hours of computation was performed on hardware of type RTX 2080 Ti (TDP of 250W). Total emissions are estimated to be 91.8 kgCO2eq of which 0 percent was directly offset. Estimations were conducted using the Machine Learning Impact calculator presented in (Lacoste et al., , 2019).
@@ Acknowledgement
Funding
No funding was received for conducting this study.
Data Availability
All the datasets used in this paper are publicly available.
Statements and Declarations
Competing Interests: The authors have no competing interests to declare that are relevant to the content of this article.